Calculate the initial concentration of Fluoroacetic acid in a solution that has the same pH as a 0.00865 M solution of HCl


1
Expert's answer
2022-04-22T03:08:04-0400

HCl is a strong acid and dissociates completely to H+ and Cl-, therefore

pH = -lg[H+] = -lgC(HCl) = -lg(0.00865) = 2.063

The equilibrium of dissociation is FCH2COOH = FCH2COO- + H+

the equilibrium constant is Ka = [H+]*[FCH2COO-]/[FCH2COOH]

C is initial concentration of the acid,

equilibrium concentrations are equal for [H+] = [FCH2COO-] = 10-2.063 = 0.00865 M,

equilibrium concentration of the acid is [FCH2COOH] = C - [H+] = C - 0.00865 M, hence:

Ka = [H+]2/(C-[H+])

10-2.59 = 0.00257 = (0.00865)2/(C-0.00865),

C = (0.00865)2/0.00257 + 0.00865 = 0.00378 M

The initial concentration is 3.78*10-3 M